OpenStudy (deadshot):

Evaluate logv3 1 over 81?

OpenStudy (deadshot):

\[\log _{3} \frac{ 1 }{ 81 }\]

OpenStudy (raden):

let will be equal x 3^x = 1/81 solve for x

OpenStudy (deadshot):

so x = 1/27?

OpenStudy (raden):

no. setting 1/81 be 1/3^4 = 3^-4

OpenStudy (deadshot):

okay

OpenStudy (raden):

now u have equation : 

OpenStudy (raden):

cancel out the 3 to both sides, exactly u will get ur answer

OpenStudy (deadshot):

so x=-4

OpenStudy (raden):

yup

OpenStudy (deadshot):

okay, thanks!

OpenStudy (raden):

you're welcome